Dog apapos owns a small fruit and vegetable store. dogg has 40 dozen tomatoes . joe expects 20% spoilage rate. the tomatoes cost 1.10 per dozen. dogg wants a 70% mar up on cost. what shoud dogg charge for each dozen tomatoes. round to the nearest cent
Answer #1

Solution:

Cost = Cost per tomato X total no. of apples = 1.10 * 40 = 44

Markup on cost = Cost * Markup rate = 44 * 70% = 30.8

Total Selling Price = 44 + 30.8 = 74.80

Loss due to Spoilage = 40*20% = 8 tomatoes

Selling price = 74.80 / (40-8) = 74.80 / 32 = $2.34 per dozen